ELECTRONIC STRUCTURE CALCULATIONS USING NONLINEAR BASIS EXPANSIONS

\underline{RON SHEPARD}^{a} and GERGELY GIDOFALVI, Chemical Sciences and Engineering Division, Argonne National Laboratory, Argonne, IL 60439.

Recent progress and applications using electronic wave functions expanded in a basis of nonlinear functions will be presented^{b}.

aThis work was supported by the Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences, and Biosciences, U.S. Department of Energy, and in part through the SciDAC project “Advanced Software for the Calculation of Thermochemical, Kinetics, and Dynamics,” under contract number DE-AC02-06CH11357.

bR. Shepard, M. Minkoff, and S. R. Brozell, Int. J. Quantum Chem. 107, 3203 (2007)